Give the formulas for and plot AFC,MC,AVC and AC if the cost function is: A. c=10+10Q, B. C=10+q2, C. C=10+10q-4q²+q3.


1
Expert's answer
2021-10-24T18:09:54-0400

Solution:

A.). C = 10 + 10Q

AFC = "\frac{TFC}{Q} = \frac{10}{Q}"


MC = "\frac{\partial TC} {\partial Q} = 10"

AVC = "\frac{VC} {Q} = \frac{10Q} {Q} = 10"


AC = "\frac{TC} {Q} = \frac{10 + 10Q} {Q} = \frac{10} {Q} + 10"

 

B.). C = 10 + q2

AFC = "\frac{TFC}{Q} = \frac{10}{Q}"


MC = "\frac{\partial TC} {\partial Q} = 2Q"


AVC = "\frac{VC} {Q} = \frac{Q^{2} } {Q} = 10"


AC = "\frac{TC} {Q} = \frac{10 + Q^{2} } {Q} = 10 + Q"


 

C.). C = 10 + 10q – 4q2 + q3

AFC = "\frac{TFC}{Q} = \frac{10}{Q}"


MC = "\frac{\partial TC} {\partial Q} = 10" "- 8Q - 3Q^{2}"


AVC = "\frac{VC} {Q} = \frac{10 + 10Q - 4Q^{2} +Q^{3} } {Q} = \frac{10} {Q} + 10 - 4Q + Q^{2}"